OpenStudy (anonymous):

Will Fan and Medal Nine and three-sevenths feet of ribbon is glued to the circumfrence of a round table what is the diameter?

OpenStudy (kamibug):

d = C / pi

OpenStudy (anonymous):

I dont know how to do it just looking at the problem i could tell you the C= 9 3/7 feet

OpenStudy (kamibug):

Right. That's the circumference. The problem is asking for the diameter, which is equal to the circumference divided by pi (22/7). :) You can solve it like this ... First change 9 and 3/7 to 66/7 Then use the "Keep, Change Flip" method to divide it by 22/7. \[d = \frac{ 66 }{ 7 } * \frac{ 7 }{ 22 }\]

OpenStudy (anonymous):

C???

OpenStudy (anonymous):

@KamiBug @AloneS

OpenStudy (kamibug):

Not quite. How did you get C ? Check your calculations again. :)

OpenStudy (anonymous):

I got C by gussing i am not very good at math. maybe B???

OpenStudy (kamibug):

B is right. Here's why ! Just multiply across the fractions... 66*7 = 462 and 7*22 = 154 Then divide 462 / 154 = 3 Get it ? :)

OpenStudy (anonymous):

Thank you!!! :)

OpenStudy (kamibug):

No problem :D Feel free to keep tagging me :P